OpenStudy (anonymous):

Your piggy bank has a total of 46 coins in it; some are dimes and some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have?

OpenStudy (anonymous):

Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770? this one two

OpenStudy (radar):

I love these problems where they can't keep track of their money or they don't know how many people went with them to the park. Which one you want to work first?

OpenStudy (anonymous):

salisbury

OpenStudy (radar):

OK, although they both are the same type of problem. You have in this case two types of people, Adults and Students (this is not to say an adult could not be a student) but in this case that is the fact. Another fact is the total "people" are 150. Lets work this with only one variable, or are you required to use two?

OpenStudy (radar):

No answer so I will use just one. Let x = number of students, then from the facts given, we can deduce that: 150-x = number of adults. That takes care of that, now oh yes the tickets. The money taken in from the tickets is as stated in the problem is a total of $4770 and would result from people buying tickets which is: 28.50x from students pockets and (150-x)45.00 from the adults pockets, Lets add those and make it equal to $4770. 28.50x +6750-45.00x=4770 dollars -16.50x=-1980 (dollars) x=1980/16.5=120 students of course that is the number of students and the adults would be 150-120 = 30 30 adults

OpenStudy (radar):

any quesrtions? note in the solution for x the dollars cancelled out leaving just a number no units. x was equal to:\[$1980\over$16.50\] for answer of 120

OpenStudy (radar):

It is rude to leave when someone is attempting to help you!

OpenStudy (radar):

monologue vs dialogue